If mean is 40, range is 10, mode is 45, variance is 144, median is 43; then find the standard deviation, find coefficient of variation, and find 50th percentile
Answer #1

Given

Mean = 40

Range = 10

Mode = 45

Variance = 144

Median = 43

1) standard deviation = \sqrt variance

Standard deviation = \sqrt 144

Standard deviation = 12

2) coefficient of variation = ( standard deviation/mean )*100 %

Coefficient of variation =( 12/40)*100 %

Coefficient of variation = 30%

3) 50th percentile is median

50th percentile = 43